OpenStudy (anonymous):

Evaluate 3(x+y) - ---- 4xy^2 when x = 2/3 and y=-2

OpenStudy (anonymous):

3(2/3 -2)/4(2/3)(4)=(-4)3/32=-3/8

jimthompson5910 (jim_thompson5910):

[ 3(x+y) ]/[ 4xy^2 ] [ 3(2/3 + (-2)) ]/[ 4(2/3)(-2)^2 ] [ 3(2/3 + (-2)) ]/[ 4(2/3)(4) ] [ 3(2/3 + (-2)) ]/[ 32/3 ] [ 3(2/3 - 2) ]/[ 32/3 ] [ 3(-4/3) ]/[ 32/3 ] [ -12/3 ]/[ 32/3 ] (-12/3)(3/32) -36/96 -3/8 So [ 3(x+y) ]/[ 4xy^2 ] = -3/8 when x = 2/3 and y = -2